OpenStudy (anonymous):

Eddie is fishing. He has 1 1/4 ounces of weight on his line, but his bait isn’t getting to the bottom of the lake. Eddie adds another1/2 -ounce weight to the line, but it's still not enough. Finally, Eddie adds a 3/4 -ounce weight and the bait sinks. How much weight does Eddie now have on his line?

OpenStudy (chriss):

2 1/2 ounces

OpenStudy (anonymous):

Trinh is catering a wedding that 112 people will attend. She is making two cakes, a chocolate cake and a carrot cake.5/8 of the guests will be served chocolate cake and the rest will receive carrot cake. How many guests will get carrot cake?

OpenStudy (chriss):

3/8(112)=42

OpenStudy (anonymous):

Two students, Raymond and Shauna, are trying to get city council members to vote for a new park along the river. 2/6 of the 24 council members say they already support the idea, but that's not enough. At least 13 council members must support the idea before the park is approved. How many more city council members must Raymond and Shauna convince?

OpenStudy (chriss):

2/6(24)=8, so 5 more council members

OpenStudy (anonymous):

Katrina is building a deck. She cuts 3 lengths of wood from a 72-inch board. The first length is 30 3/4 inches. The second length is 12 1/4 inches. And the third length is 16 1/4 inches. How much of the original board remains after Katrina makes her cuts?

OpenStudy (chriss):

12 3/4 inches

OpenStudy (anonymous):

Find the product. Select the answer that is in simplest form. 3/4×3/5=

OpenStudy (chriss):

9/20

OpenStudy (anonymous):

its 5